The raw data from grain-size test is shown below. What is the value of Cc and Cu?

Sieve #

Sieve Size (mm)

Mass Retained (g)

¾ “

19.00

773

# 4

4.75

3316.1

# 10

2.00

298

# 30

0.595

616.1

# 50

0.297

403.9

# 100

0.150

202.4

# 200

0.075

52.2

Pan

14.4
Answer #1

100 10 0.1 0.01 Sieve Size (mm)D10 = 0.5 mm

D30 = 5 mm

D60 = 10.6 mm

Cu = D60/D10 = 10.6/0.5 = 21.2

Cc = D302/(D60*D10) = 52/(10.6 * 0.5) = 4.72